Reduce the quadratic form Q=x1²+2x2x3 to canonical form and hence find its nature, rank, index and signature.
1
Expert's answer
2020-11-25T16:58:52-0500

"Q=x^2_1+2x_2x_3"

Let

"x_1 =y_1\\
x_2=y_2-y_3\\
x_3=y_2+y_3"

"Q=y^2_1+2(y_2-y_3)(y_2+y_3)=\\
=y^2_1+2y^2_2-2y^2_3"

canonical form is "Q=y^2_1+2y^2_2-2y^2_3"

rank=3 (№ of non-zero eigen values)

index= 2 (№ of positive eigen values)

signature=2-1=1 (difference betwen № of positive and negative eigen values)

nature: indefinite if some of the eigen values of Q are + ve and others – ve.